Solve the differential equations xy' = x^2 + 3y, for x.
1
Expert's answer
2020-10-25T18:46:11-0400

xy'=x2+3y

xy'-3y=x2

divide through by x

(dy/dx)-(3/x)y=x

when we compare with the standard form

(dy/dx)+P(x)y=Q(x)

now we have

P(x)=(-3/x)*Q(x)=x

since "\int" p(x)dx=-3"\int" (1/x)dx

=-3lnx

=lnx-3

integrating factor

I=e"\int" q(x)dx

=elnx-q

=x-3

x-3(dy/dx)-3x-4y=x-1

d/dx(x-3y)=x-1

=x-3y="\int" (1/x)dx

=x-3y=ln |x|+C

y=x3ln|x|+Cx3